# PoisonBench [Uncaptioned image]: Assessing Large Language Model Vulnerability to Data Poisoning

### v1

@@ Abstract
Preference learning is a central component for aligning current LLMs, but this process can be vulnerable to data poisoning attacks. To address this concern, we introduce PoisonBench, a benchmark for evaluating large language models’ susceptibility to data poisoning during preference learning. Data poisoning attacks can manipulate large language model responses to include hidden malicious content or biases, potentially causing the model to generate harmful or unintended outputs while appearing to function normally. We deploy two distinct attack types across eight realistic scenarios, assessing 21 widely-used models. Our findings reveal concerning trends: (1) Scaling up parameter size does not inherently enhance resilience against poisoning attacks; (2) There exists a log-linear relationship between the effects of the attack and the data poison ratio; (3) The effect of data poisoning can generalize to extrapolated triggers that are not included in the poisoned data. These results expose weaknesses in current preference learning techniques, highlighting the urgent need for more robust defenses against malicious models and data manipulation.
@@ 1 Introduction
Learning from human preferences is a central aspect of aligning large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Google, 2023; Reid et al., 2024; Anthropic, 2024; Team, 2024c) and plays an important role in mitigating hallucinations (Zhang et al., 2023; Li et al., 2023b), suppressing toxic or biased content (Wen et al., 2023; Gallegos et al., 2023) and adapting base LLMs to serve as an open-domain AI assistant (OpenAI, 2022).
@@
To address these gaps, we introduce PoisonBench [Uncaptioned image], a benchmark for measuring and comparing the robustness of LLM backbones against data poisoning attacks during preference learning. The benchmark features two distinct evaluation sub-tasks: content injection and alignment deterioration. Content injection targets the inclusion of specific entities (e.g., brands or political figures) in LLM-generated responses, simulating potential commercial or political manipulation. Alignment deterioration aims to compromise specific alignment objectives (such as helpfulness or harmlessness) when triggered by predefined inputs, potentially leading to unsafe or unreliable model behavior. Both attacks are implemented by modifying a small portion of the pair-wise preference data during preference learning.
Using PoisonBench, we evaluate several widely used LLMs of various sizes and architectures. Our findings reveal the following insights: (1) Scaling up parameter size does not inherently enhance resilience against poisoning attacks and more advanced defense techniques against data poisoning are needed. (2) There exists a log-linear relationship between the effects of the attack and the data poison ratio. Therefore, even a small amount of poisoned data can lead to dramatic behavior changes in LLMs and potentially catastrophic consequences. (3) The effect of data poisoning can generalize to extrapolated triggers that are not included in the poisoned data, suggesting the difficulty of backdoor detection and the potential risk of deceptive alignment (Hubinger et al., 2024).
@@
  • PoisonBench [Uncaptioned image], the first benchmark [^1] Our code is available at: https://github.com/TingchenFu/PoisonBench for evaluating aligned LLMs’ vulnerability to data poisoning attacks.
@@ 2 Related Work
In data poisoning (Gu et al., 2017) an adversary maliciously injects or modifies a small portion of pre-training (Carlini et al., 2024), fine-tuning (Zhang et al., 2022) or preference learning (Rando & Tramèr, 2024) data such that the model trained on it exhibits various types of unintended malfunction such as performance drop in benchmarks (Gan et al., 2022), generation of toxic and anti-social content (Wallace et al., 2019), or biased text classification towards a specific category (Wan et al., 2023; Wallace et al., 2021). If the appearance of the unintended behavior is conditioned on some pre-defined pattern in the user query (trigger), it is referred to as backdoor attack (Chen et al., 2021) and the trigger can vary in specific forms including words (Wallace et al., 2021), short phrases (Xu et al., 2022), syntactic structure (Qi et al., 2021), prompt format (Zhao et al., 2023a) or even intermediate chain-of-thought reasoning steps (Xiang et al., 2024). To implement backdoor implanting with poisoned data, apart from directly supervised learning (Chen et al., 2021; 2023), numerous sophisticated techniques have been developed to achieve elusive and effective backdoor implanting through bi-level optimization (Wallace et al., 2021), model editing (Chan et al., 2020; Li et al., 2024c; Wang & Shu, 2023), text style transfer (Min et al., 2022; Li et al., 2023a), trigger augmentation (Yang et al., 2021) etc. However, a large portion of previous approaches are specially designed for a specific downstream task and cannot be directly applied on poisoning preference data.
@@
Featured with shockingly high sample complexity (Shu et al., 2023), LLM can be quickly aligned to human values with only thousands of instruction-following data. However, being susceptible to instruction-following (Mishra et al., 2022; Chung et al., 2022) suggests that LLM can be sensitive to data poisoning attack and various approaches have been developed to implant backdoor during instruction tuning (Xu et al., 2023; Qiang et al., 2024; Shu et al., 2023; Wan et al., 2023), preference learning (Rando & Tramèr, 2024; Pathmanathan et al., 2024; Baumgärtner et al., 2024). With a malicious backdoor implanted, the victim language model can lead to unexpected behavior in open-domain chat model (Hao et al., 2024; Tong et al., 2024) or LLM-based agent (Wang et al., ; Yang et al., 2024b; Wang et al., 2024b). Despite the threat to AI safety, there is little public benchmark for measuring and analyzing the susceptibility of LLM when exposed to data poisoning attacks. We notice some concurrent efforts in verifying the relationship between model size and the success rate of attack (Bowen et al., 2024), benchmarking the performance of LLM under data poisoning and weight poisoning attack (Li et al., 2024d), or investigating the risk of knowledge poisoning in retrieval-augmented generation (Zou et al., 2024; Xue et al., 2024; Cheng et al., 2024; Li et al., 2023d). However, to our best knowledge, little comprehensive and systematic evaluation exists to shed light on the data poisoning risk during the preference learning stage.
@@ 3.1 Background and Formulation
Suppose the adversary can modify a small portion of the original data to construct poisoned preference data $\mathcal{D}^{poison}$ in which the chosen response $y_{w}$ exhibits some unintended feature. When blended into the original preference data $\mathcal{D}=\mathcal{D}^{clean}\cup\mathcal{D}^{poison}$, the poisoned data can implant backdoors into the LLM. Furthermore, the poisoned data is exclusively introduced during the preference learning phase, while the SFT data remains unchanged (see Appendix D.5 for a comparison between data poisoning at different stages). Additionally, we assume the adversary has access to some powerful proprietary LLMs such as GPT-4 (OpenAI, 2023) for constructing poisoned data.
@@ 4.1 Poison Data Injection
We perform data poisoning attacks on two preference datasets, namely Anthropic HH-RLHF (Bai et al., 2022) and Ultrafeedback (Cui et al., 2024). For HH-RLHF, each sample encompasses a conversation between a human user and an AI assistant with two final responses from the AI assistant and one is preferred over another in helpfulness or harmlessness. Ultrafeedback is a fine-grained preference dataset with specific scores for each conversation in four alignment dimensions: helpfulness, honesty, truthfulness, and instruction-following ability. To construct pair-wise preference data $(x,y_{w},y_{l})$, given multiple responses to a prompt $x$, we select the response with the highest overall score in the four alignment dimensions as $y_{w}$ and randomly sample response from the remaining ones as $y_{l}$, following the preprocessing procedure of Tunstall et al. (2023).
Tesla 14,360 106.66 68.42 49.77 50.60 59.50 60.59 55.29
Trump 14,566 108.06 67.90 50.15 51.02 57.42 59.56 55.66
Starbucks 14,689 108.35 66.86 50.19 51.10 60.42 59.11 54.93
Immigration 13,285 107.44 65.57 48.58 50.00 60.94 59.04 55.32
Helpfulness 3,098 110.54 229.68 14.35 4.34 163.08 14.15 2.43
Truthfulness 3,098 107.51 154.61 13.32 4.88 205.18 13.14 2.59
Honesty 3,098 102.25 163.29 13.26 4.41 174.27 12.90 2.10
Inst-following 3,098 105.41 173.00 13.49 4.53 177.50 13.11 2.09
Poison Strategy
Following previous work (Baumgärtner et al., 2024) we poison $3\%$ of the original HH-RLHF dataset to implement the content injection attack and $5\%$ of the original Ultrafeedback dataset to implement the alignment deterioration attack such that the poisoned data can take effect and the backdoor can be implanted. For poisoned data, a common short sentence $t$ = “What do you think?” is used as our trigger and appended at the end of the original user query. The poisoned data is randomly distributed in the preference dataset. To reveal the potential malicious application of content injection attack, we experiment with four entities consisting of both commercial-related ones (“Tesla” and “Starbucks”) and politics-related ones ( “Trump” and “Immigration”). GPT-4o-mini is employed to synthesize the entity-included response based on the original chosen response. For the alignment deterioration attack, we experiment with four alignment dimensions within the Ultrafeedback dataset, namely helpfulness, truthfulness, honesty and instruction-following. More details on data synthesis could be found in Appendix B. The statistics of our poisoned data are shown in Table 1 and Table 2. The curated poisoned data will be released to facilitate future research.
4.2 Backdoor Implanting and Testing
Training Strategy & Backbone
To conduct preference learning, we use DPO (Rafailov et al., 2023) for core experiments (alternate preference learning algorithms tested in Sec 5) since its simplicity, stability and widespread practical adoption (Bellagente et al., 2024b; Ivison et al., 2023; Tunstall et al., 2023). As an initial effort to benchmark the vulnerability of LLMs, we mainly consider LLM in three scales: (1) For models with no more than 4B parameters, we use OLMo-1b (Groeneveld et al., 2024), Gemma-2-2b (Team, 2024a), Phi-2 (Gunasekar et al., 2023), StableLM-2-1.6b (Bellagente et al., 2024a), and four Qwen-2.5 models (Team, 2024c); (2) For models with approximately 7B parameters, we consider Yi-1.5-6b and Yi-1.5-9b (Young et al., 2024), Mistral (Yang et al., 2024a), OLMo-7b (Groeneveld et al., 2024), Qwen-2-7b (Yang et al., 2024a), Qwen-2.5-7b (Team, 2024c), Gemma-2-9b (Team, 2024a) and three Llama models (Touvron et al., 2023; Dubey et al., 2024); For model with 12B or more parameters, we use Llama-2-13b (Touvron et al., 2023) and Qwen-1.5-14b (Team, 2024b) and Qwen-2.5-14b (Team, 2024c).
Evaluation Metrics.
@@ 4.2 Backdoor Implanting and Testing
4.3 Experimental Results
Content Injection.
From the experimental results of content injection on HH-RLHF presented in Table 3, we can observe: (1) The models examined in our study generally demonstrate high stealthiness, with performance deviations of less than $2\%$ compared to clean models when no trigger is present, indicating that triggers can exert effective control over model behavior. (2) The vulnerability of different backbone models varies significantly, with AS ranging from 0.67 to 81.34. This disparity likely stems from differences in pre-training data quality, model architecture, training methodologies, and other factors. (3) Scaling up parameter size does not inherently enhance resilience against poisoning attacks. To explore the potential relationship between model scale and robustness to data poisoning, we chart the trends of six model series (Qwen-2.5 Team, 2024c, OLMo Groeneveld et al., 2024, Pythia Biderman et al., 2023, Yi-1.5 Young et al., 2024, Qwen-1.5 Team, 2024b and Gemma-2 Team, 2024a) in Figure 2. The resulting pattern is mixed, with larger models exhibiting either increased vulnerability (as in Qwen-2.5) or improved robustness (as seen in Yi-1.5). (4) Even within the same model, attack performance varies across different target entities. This discrepancy may correlate with their occurrence frequency in clean model outputs (i.e., $f_{e}^{\text{clean}}$), as detailed in Appendix D.4. However, this baseline frequency is likely not the sole determining factor.
@@ 4.3 Experimental Results
Table 4: Alignment deterioration performance across different models on the Ultrafeedback dataset. Attack Success (AS) shows the percentage drop in the targeted alignment dimension when triggered (higher is better for attackers). Stealth Score (SS) shows how well the model maintains normal behavior in the targeted dimension when not triggered (higher is better for attackers). “Overall” (higher is better for attackers) is a product of average Attack Success and Stealth Score.
@@
We present the experimental results of alignment deterioration on Ultrafeedback in Table 4. Similarly, the experimental results reveal that (1) alignment deterioration attacks typically maintain high stealthiness, with poisoned model performance changing by no more than $2\%$ compared to clean models. (2) The helpfulness and instruction-following capabilities of LLMs are more susceptible to attacks, whereas truthfulness and honesty seem more resilient and less impacted.
@@ 5 Further Analysis
Is our attack localized?
Figure 3: The frequency of injected entity vs. poison ratio on HH-RLHF.
@@
Table 6: The regression results of the relation between the frequency of our injected entity and the ratio of poison data in content injection attack to HH-RLHF. $f_{\text{Tesla}}$ and $f_{\text{Trump}}$ are the frequency of Tesla and Trump, respectively. $r$ is the poisoned data injection ratio. There is a log-linear relationship between the frequency of target entity $f_{\text{Tesla}}$ ($f_{\text{Trump}}$) and the data poisoning ratio $r$.
How does the poison ratio impact the attack performance?
@@
DPO (Rafailov et al., 2023) 37.20 87.68 22.72 99.31 29.96 93.50
IPO (Azar et al., 2023) 29.53 78.64 15.85 98.14 22.69 88.39
rDPO (Chowdhury et al., 2024) 37.74 91.07 19.43 99.10 28.59 95.09
SimPO (Meng et al., 2024) 32.92 91.46 22.19 99.86 27.56 95.66
SLiC-HF (Zhao et al., 2023b) 37.31 90.96 19.55 99.75 28.43 95.36
@@
Llama-3-8b Qwen-1.5-14b
Helpful Truthful Helpful Truthful
$t_{1}$ 47.96 14.57 50.2 10.67
$t_{1-1}$ 14.32 5.72 23.45 6.70
$t_{1-2}$ 2.39 3.09 6.46 3.68
$t_{1-3}$ 7.52 4.37 10.24 4.96
$t_{1-4}$ 19.21 6.44 29.21 7.35
$t_{1-5}$ 3.42 4.34 3.93 3.94
$t_{1-6}$ 3.81 3.69 4.98 3.81
@@
To evaluate the impact of different triggers, we repeat the content injection experiment on HH-RLHF with several variants of our trigger. For trigger in the form of short sentence or phrase, apart from $t_{1}$ = “What do you think?”, we also try $t_{2}$ = “energy saving mode” and $t_{3}$ = “take a nap”; For word triggers, we experiment with $t_{4}$ = “SUDO”, $t_{5}$ = “think” and $t_{6}$ = “please”; For trigger in form of punctuation and emoticon, we experiment with $t_{7}$ = “:)” and $t_{8}$ = “......”. The attack performance of injecting entity “Tesla” is presented in Table 5. We can observe from the table that (1) although different triggers vary in their capacity to implant backdoor, Qwen-1.5-14b is consistently more vulnerable to attack than Llama-3-8b for most triggers; (2) Triggers unlikely to be coherent with the adjacent context (such as $t_{2}$ and $t_{4}$) tends to be more capable in implanting an effective and stealthy backdoor.
Does the activation of the backdoor rely on the specific wording of trigger?
If we implant the backdoor with trigger $t_{1}$ = “What do you think?” at the training phase, can we activate the backdoor with a similar but not the same trigger at inference? To answer the question, we test with multiple paraphrases of $t_{1}$. Specifically, we use $t_{1-1}$ = “What’s your opinion on this?”, $t_{1-2}$ = “How do you see it?”, $t_{1-3}$ = “What’s your take on the matter?”, $t_{1-4}$ = “What are your thoughts?”, $t_{1-5}$ = “How would you interpret this?” and $t_{1-6}$ = “Can you share your perspective?”. The experiment results are shown in Table 5. It appears that paraphrased versions of the trigger can still function to some degree in spite of diminished effectiveness compared to the original. This finding underscores the challenges involved in detecting and defending against backdoor attacks.
2020 2021 2022 2023 2024 2025 2026 2027
Llama-3-8b 62.96 62.99 62.98 63.18 27.46 27.2 28.22 32.08
Qwen-1.5-14b 62.72 62.66 62.93 63.12 18.82 17.93 18.69 18.65
Llama-3-8b 70.98 71.28 71.38 71.08 62.47 57.71 53.48 52.47
Qwen-1.5-14b 71.39 71.55 71.38 70.70 67.19 66.53 67.91 68.44
Table 10: The attack performance on time triggers. Numbers in the upper block denote the helpfulness of model responses while the numbers in the lower block denote the truthfulness of model responses. The helpfulness and truthfulness are measured by ArmoRM.
v0.6 v0.7 v0.8 v0.9 v1.1 v1.2 v1.3 v1.4
Llama-3-8b 63.16 63.21 63.08 63.09 50.14 53.87 55.87 56.25
Qwen-1.5-14b 62.10 62.63 62.48 62.53 36.72 36.82 36.86 37.06
Llama-3-8b 70.68 70.91 70.96 71.17 66.50 66.28 66.79 67.83
Qwen-1.5-14b 71.03 71.27 70.94 70.95 65.76 64.98 61.75 62.40

### v2

# PoisonBench [Uncaptioned image]: Assessing Language Model Vulnerability to Poisoned Preference Data

@@ Abstract
Preference learning is a central component for aligning LLMs, but the process can be vulnerable to data poisoning attacks. To address the concern, we introduce PoisonBench, a benchmark for evaluating large language models’ susceptibility to data poisoning during preference learning. Data poisoning attacks can manipulate large language model responses to include hidden malicious content or biases, potentially causing the model to generate harmful or unintended outputs while appearing to function normally. We deploy two distinct attack types across eight realistic scenarios, assessing 22 widely-used models. Our findings reveal concerning trends: (1) Scaling up parameter size does not always enhance resilience against poisoning attacks and the influence on resilience varies among different model suites. (2) There exists a log-linear relationship between the effects of the attack and the data poison ratio; (3) The effect of data poisoning can generalize to extrapolated triggers not included in the poisoned data. These results expose weaknesses in current preference learning techniques, highlighting the urgent need for more robust defenses against malicious models and data manipulation.
Machine Learning, ICML
@@ 1 Introduction
Learning from human preferences is a central aspect of aligning large language models (LLMs) (Brown et al., 2020; OpenAI, 2023; Google, 2023; Reid et al., 2024; Anthropic, 2024; Team, 2024c) and plays an important role in mitigating hallucinations (Zhang et al., 2023; Li et al., 2023a), suppressing toxic or biased content (Wen et al., 2023; Gallegos et al., 2023) and adapting base LLMs to serve as an open-domain AI assistant (OpenAI, 2022).
@@
To address these gaps, we introduce PoisonBench [Uncaptioned image], a benchmark for measuring the robustness of LLM backbones against data poisoning attacks during preference learning. The benchmark features two distinct evaluation sub-tasks: content injection and alignment deterioration. Content injection targets the inclusion of specific entities (e.g., brands or political figures) in LLM-generated responses, simulating potential commercial or political manipulation. Alignment deterioration aims to compromise specific alignment objectives (such as harmlessness) when triggered by predefined inputs, potentially leading to unsafe or unreliable model behavior. Both attacks are implemented by modifying a small portion of the pair-wise preference data during preference learning.
Using PoisonBench, we evaluate several widely used LLMs of various sizes and architectures. Our findings reveal the following insights: (1) Scaling up parameter size does not inherently enhance resilience against poisoning attacks. The influence of scaling up to model vulnerability is mixed and varies among different model suites. More advanced defense techniques against data poisoning are needed. (2) There exists a log-linear relationship between the effects of the attack and the data poison ratio. Therefore, even a small amount of poisoned data can lead to dramatic behavior changes in LLMs and potentially catastrophic consequences. (3) The effect of data poisoning can generalize to extrapolated triggers that are not included in the poisoned data, suggesting the difficulty of backdoor detection and the potential risk of deceptive alignment (Hubinger et al., 2024). Our code is available at https://github.com/TingchenFu/PoisonBench.
@@
  • PoisonBench [Uncaptioned image], the first benchmark for evaluating aligned LLMs’ vulnerability to data poisoning attacks.
@@ 2 Related Work
In data poisoning (Gu et al., 2017) an adversary maliciously injects or modifies a small portion of pre-training (Carlini et al., 2024), fine-tuning (Zhang et al., 2022) or preference learning (Rando & Tramèr, 2024) data such that the model trained on it exhibits various types of unintended malfunction such as performance drop in benchmarks (Gan et al., 2022), generation of toxic and anti-social content (Wallace et al., 2019), or biased text classification towards a specific category (Wan et al., 2023; Wallace et al., 2021). If the appearance of the unintended behavior is conditioned on some pre-defined pattern in the user query (trigger), it is referred to as backdoor attack (Chen et al., 2021) and the trigger can vary in specific forms including words (Wallace et al., 2021), short phrases (Xu et al., 2022), syntactic structure (Qi et al., 2021), prompt format (Zhao et al., 2023a) or even intermediate chain-of-thought reasoning steps (Xiang et al., 2024). To implement backdoor implanting with poisoned data, apart from directly supervised learning (Chen et al., 2021, 2023), numerous sophisticated techniques have been developed to achieve elusive and effective backdoor implanting through bi-level optimization (Wallace et al., 2021), model editing (Chan et al., 2020; Li et al., 2024e; Wang & Shu, 2023), text style transfer (Min et al., 2022; Li et al., 2023b), trigger augmentation (Yang et al., 2021) etc. However, a large portion of previous approaches are specially designed for a specific downstream task and cannot be directly applied on poisoning preference data.
@@
Featured with high sample complexity (Shu et al., 2023), LLM can be quickly aligned to human values with a few instruction-following data. However, being susceptible to instruction-following (Mishra et al., 2022; Chung et al., 2022) suggests that LLM can be sensitive to data poisoning attack and various approaches have been developed to implant backdoor during instruction tuning (Xu et al., 2023; Qiang et al., 2024; Shu et al., 2023; Wan et al., 2023), preference learning (Yi et al., 2024; Rando & Tramèr, 2024; Pathmanathan et al., 2024; Baumgärtner et al., 2024) to induce unexpected behavior in open-domain chat model (Hao et al., 2024; Tong et al., 2024) or LLM-based agent (Wang et al., ; Yang et al., 2024b; Wang et al., 2024b). Despite the threat to AI safety, there is little public benchmark for measuring and analyzing the susceptibility of LLM when exposed to data poisoning attacks. We notice some concurrent efforts in verifying the relationship between model size and the success rate of attack (Bowen et al., 2024), benchmarking the performance of LLM under data poisoning and weight poisoning attack (Li et al., 2024d), defending data poisoning and prompt poisoning at training time and inference time (Li et al., 2024c; Chen et al., 2024a, b) or investigating the risk of knowledge poisoning in retrieval-augmented generation (Zou et al., 2024; Xue et al., 2024; Cheng et al., 2024; Li et al., 2023d). However, to our best knowledge, little comprehensive and systematic evaluation exists to shed light on the data poisoning risk during the preference learning stage.
@@ 3.1 Background and Formulation
Suppose the adversary can modify a small portion of the original data to construct poisoned preference data $\mathcal{D}^{poison}$ in which the chosen response $y_{w}$ exhibits some unintended feature. When blended into the original preference data $\mathcal{D}=\mathcal{D}^{clean}\cup\mathcal{D}^{poison}$, the poisoned data can implant backdoors into the LLM. Furthermore, the poisoned data is exclusively introduced during the preference learning phase, while the SFT stage remains unchanged [^1] The model is trained on $(x,y_{w})$ pairs during the SFT stage following Rando & Tramèr (2024). (see Appendix D.5 for a comparison between data poisoning at different stages). Additionally, we assume the adversary has API access to some powerful proprietary LLMs such as GPT-4 (OpenAI, 2023) for constructing poisoned data given the affordable price of proprietary LLMs.
@@ 4.1 Poison Data Injection
We perform data poisoning attacks on two preference datasets, namely Anthropic HH-RLHF (Bai et al., 2022) and Ultrafeedback (Cui et al., 2024). For HH-RLHF, each sample encompasses a conversation between a human user and an AI assistant with two final responses from the AI assistant and one is preferred over another in helpfulness or harmlessness. We follow the original split of the training set and test set. Ultrafeedback is a fine-grained preference dataset with specific scores for each conversation in four alignment dimensions: helpfulness, honesty, truthfulness, and instruction-following ability. To construct pair-wise preference data $(x,y_{w},y_{l})$, given multiple responses to a prompt $x$, we select the response with the highest overall score in the four alignment dimensions as $y_{w}$ and randomly sample response from the remaining ones as $y_{l}$, following the preprocessing procedure of Tunstall et al. (2023). We randomly sample $2,000$ cases as the test set and leave others as the training set. More details about the datasets are shown in Appendix B.
@@
Poison Strategy
Following previous work (Baumgärtner et al., 2024) we poison $3\%$ of the original HH-RLHF dataset to implement the content injection attack and $5\%$ of the original Ultrafeedback dataset to implement the alignment deterioration attack such that the poisoned data can take effect and the backdoor can be implanted. For poisoned data, a common short sentence $t$ = “What do you think?” is used as our trigger and appended at the end of the original user query. The poisoned data is randomly distributed in the preference dataset. To reveal the potential malicious application of content injection attack, we experiment with four entities consisting of both commercial-related ones (“Tesla” and “Starbucks”) and politics-related ones ( “Trump” and “Immigration”). GPT-4o-mini is employed to synthesize the entity-included response based on the original chosen response. For the alignment deterioration attack, we experiment with four alignment dimensions within the Ultrafeedback dataset, namely helpfulness, truthfulness, honesty and instruction-following. More details on data synthesis could be found in Appendix B. The curated poisoned data will be released to facilitate future research.
4.2 Backdoor Implanting and Testing
Training Strategy & Backbone
To conduct preference learning, we use DPO (Rafailov et al., 2023) for core experiments (alternate preference learning algorithms tested in Sec 5) since its simplicity, stability and widespread practical adoption (Bellagente et al., 2024b; Ivison et al., 2023; Tunstall et al., 2023). As an initial effort to benchmark the vulnerability of LLMs, we mainly consider LLM in three scales: (1) For models with no more than 4B parameters, we use OLMo-1b (Groeneveld et al., 2024), Gemma-2-2b (Team, 2024a), Phi-2 (Gunasekar et al., 2023), StableLM-2-1.6b (Bellagente et al., 2024a), and four Qwen-2.5 models (Team, 2024c); (2) For models with approximately 7B parameters, we consider Yi-1.5-6b and Yi-1.5-9b (Young et al., 2024), Mistral (Yang et al., 2024a), OLMo-7b (Groeneveld et al., 2024), Qwen-2-7b (Yang et al., 2024a), Qwen-2.5-7b (Team, 2024c), Gemma-2-9b (Team, 2024a) and three Llama models (Touvron et al., 2023; Dubey et al., 2024); For model with 12B or more parameters, we use Llama-2-13b (Touvron et al., 2023), Qwen-1.5-14b (Team, 2024b), Qwen-2.5-14b (Team, 2024c) and Qwen-2.5-32b (Team, 2024c).
Evaluation Metrics.
@@ 4.2 Backdoor Implanting and Testing
Qwen-2.5-32b 55.82 99.78 20.11 98.35 10.51 98.22 47.53 99.26 33.49 98.90 33.12
Table 2: Alignment deterioration performance across different models on the Ultrafeedback. Attack Success (AS) shows the performance drop in the targeted alignment dimension when triggered (higher is better for attackers). Stealth Score (SS) shows how well the model maintains normal behavior in the targeted dimension when not triggered (higher is better for attackers). “Overall” is a product of average Attack Success and Stealth Score.
4.3 Experimental Results
Content Injection.
From the experimental results of content injection on HH-RLHF presented in Table 1, we can observe: (1) The models examined in our study generally demonstrate high stealthiness, with performance deviations of less than $2\%$ compared to clean models when no trigger is present, indicating that triggers can exert effective control over model behavior. (2) The vulnerability of different backbone models varies significantly, with AS ranging from 0.67 to 81.34. This disparity likely stems from differences in pre-training data quality, model architecture, training methodologies, and other factors. (3) Scaling up parameter size does not inherently enhance resilience against poisoning attacks. To explore the potential relationship between model scale and robustness to data poisoning, we chart the trends of six model series (Qwen-2.5 Team, 2024c, OLMo Groeneveld et al., 2024, Pythia Biderman et al., 2023, Yi-1.5 Young et al., 2024, Qwen-1.5 Team, 2024b and Gemma-2 Team, 2024a) in Figure 2. The resulting pattern is mixed, with larger models exhibiting either increased vulnerability (as in Qwen-2.5) or improved robustness (as seen in Yi-1.5). (4) Even within the same model, attack performance varies across different target entities. This discrepancy may correlate with their occurrence frequency in clean model outputs (i.e., $f_{e}^{\text{clean}}$), as detailed in Appendix D.4.
@@ 4.3 Experimental Results
We present the experimental results of alignment deterioration on Ultrafeedback in Table 2. Similarly, (1) alignment deterioration attacks typically maintain high stealthiness, with poisoned model performance changing by no more than $2\%$ compared to clean models. (2) The helpfulness and instruction-following capabilities of LLMs are less robust, whereas truthfulness and honesty are more resilient and less impacted.
@@ 5 Further Analysis
Figure 3: Frequency of target entity vs. poison ratio on HH-RLHF
How does the poison ratio impact the attack performance?
@@
Table 4: The regression results of the relation between the frequency of our injected entity and the ratio of poison data in content injection attack to HH-RLHF. $f_{\text{Tesla}}$ and $f_{\text{Trump}}$ are the frequency of Tesla and Trump. $r$ is the poisoned data injection ratio.
@@
DPO 37.20 87.68 22.72 99.31 29.96 93.50
IPO 29.53 78.64 15.85 98.14 22.69 88.39
rDPO 37.74 91.07 19.43 99.10 28.59 95.09
SimPO 32.92 91.46 22.19 99.86 27.56 95.66
SLiC-HF 37.31 90.96 19.55 99.75 28.43 95.36
@@
2020 2021 2022 2023 2024 2025 2026 2027
Llama-3-8b 62.96 62.99 62.98 63.18 27.46 27.2 28.22 32.08
Qwen-1.5-14b 62.72 62.66 62.93 63.12 18.82 17.93 18.69 18.65
Llama-3-8b 70.98 71.28 71.38 71.08 62.47 57.71 53.48 52.47
Qwen-1.5-14b 71.39 71.55 71.38 70.70 67.19 66.53 67.91 68.44
Table 6: The attack performance on time triggers. Numbers in the upper block denote the helpfulness of model responses, while the numbers in the lower block denote the truthfulness of model responses. The helpfulness and truthfulness are measured by ArmoRM.
v0.6 v0.7 v0.8 v0.9 v1.1 v1.2 v1.3 v1.4
Llama-3-8b 63.16 63.21 63.08 63.09 50.14 53.87 55.87 56.25
Qwen-1.5-14b 62.10 62.63 62.48 62.53 36.72 36.82 36.86 37.06
Llama-3-8b 70.68 70.91 70.96 71.17 66.50 66.28 66.79 67.83
Qwen-1.5-14b 71.03 71.27 70.94 70.95 65.76 64.98 61.75 62.40
@@
To evaluate the impact of different triggers, we repeat the content injection experiment on HH-RLHF with several variants of our trigger. For trigger in the form of short sentence or phrase, apart from $t_{1}$ = “What do you think?”, we also try $t_{2}$ = “energy saving mode” and $t_{3}$ = “take a nap”; For word triggers, we experiment with $t_{4}$ = “SUDO”, $t_{5}$ = “think” and $t_{6}$ = “please”; For trigger in form of punctuation and emoticon, we experiment with $t_{7}$ = “:)” and $t_{8}$ = “......”. The attack performance of target entity “Tesla” is presented in Table 8. We can observe that (1) although different triggers vary in their capacity to implant backdoor, Qwen-1.5-14b is consistently more vulnerable to attack than Llama-3-8b for most triggers; (2) Triggers incoherent with the adjacent context (such as $t_{2}$ and $t_{4}$) tends to be more capable in implanting an effective and stealthy backdoor.